1. A tool rental company charges a basic fee to rent tools. An additional amount is added to this fee for each day that the tool is rented.
It costs $24 to rent a particular tool for 2 days. For 5 days, the rental cost is $45.
How much would it cost to rent this tool for 8 days?



Answer by Abbey(339)

Let x = the basic fee
let y = the additional amount charged
basic fee + number of days*charge = total cost
x+5y=45
x+2y=24
subtract the second equation from the first:
3y=21
divide both sides by 3
y=$7
put y back into either equation and solve for x:
x+5(7)=45
x+35=45
x=10
verify with other equation:
x+2y=24
10+2(7)=24
10+14=24, so this is correct.
To find out how much it costs for 8 days,
basic charge + 8days(cost per day)=total
10+8(7)= total
10+56=66
$66 for 7 day rental